### v1

# ILID: Native Script Language Identification for Indian Languages

Abstract
Language Identification task is a crucial fundamental step in NLP. Often it serves as a pre-processing step for widely used NLP applications such as multilingual machine translation, information retrieval, question and answering, and text summarization. The core challenge of language identification lies in distinguishing languages in noisy, short, and code-mixed environments. This becomes even harder in case of diverse Indian languages that exhibit lexical and phonetic similarities, but have distinct differences. Many Indian languages share the same script making the task even more challenging. In this paper, we release a dataset of 230K sentences consisting of English and all 22 official Indian languages labeled with their language identifiers where data in most languages are newly created. We also develop and release robust baseline models using state-of-the-art approaches in machine learning and deep learning that can aid the research in this field. Our baseline models are comparable to the state-of-the-art models for the language identification task.
1 Introduction
India is a linguistically diverse country. Although there are more than 1000 languages in the Indian subcontinent, the digital divide among many Indian languages is enormous. Barring a few languages, most of the languages suffer from resource scarcity and a simple task such as Language IDentification (LID) remains a challenging task Caswell et al. (2020) for them. Available LID tools such as Lui and Baldwin (2011), Joulin et al. (2016a), and Team et al. (2022) perform poorly on Indian languages and they do not cover many of them. With the increase in mobile and internet users in India, the need for providing services in native Indian languages is more pressing than ever. Hence, it becomes essential to create robust LID tools to cater to the Indian users. As the volume and variety of data keeps growing, we need to include data from different domains while creating sound benchmarks. Keeping this motivation in mind, we set out to create an Indian Language IDentification (ILID) benchmark including English and 22 official languages consisting of a total 230K sentences labeled with their language markers. The main contribution of our paper is two folds:
  • We create ILID dataset consisting of newly created datasets in 13 languages and curated datasets in rest of the languages.
  • We develop ILID baseline models exploring different machine learning and deep learning techniques.
2 ILID Dataset Creation
Indian Language IDentification (ILID) dataset is created using two approaches. We include English and all 22 official Indian languages  [^1] https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India widely used in India. The first approach utilizes web scraping for the languages in which the digital presence is significant (details are given in the Appendix). For each of these languages, we collect 10,000 different sentences from diverse sources such as government websites, newspapers, books, and other public materials ensuring varying degrees of linguistic complexity. The second approach samples sentences from an existing massive monolingual and parallel corpora for Indian languages Mujadia and Sharma (2024). The details of the dataset are presented in Table 1. The data in each language is split into 80:10:10 ratio to create train, dev, and test sets.
Language #Train #Dev #Test #Total
Assamese (asm) 8000 1000 1000 10000
Bengali (ben) 8000 1000 1000 10000
Bodo (brx) 8000 1000 1000 10000
Dogri (doi) 8000 1000 1000 10000
Konkani (gom) 8000 1000 1000 10000
Gujarati (guj) 8000 1000 1000 10000
Hindi (hin) 8000 1000 1000 10000
Kannada (kan) 8000 1000 1000 10000
Kashmiri (kas) 8000 1000 1000 10000
Maithili (mai) 8000 1000 1000 10000
Malayalam (mal) 8000 1000 1000 10000
Marathi (mar) 8000 1000 1000 10000
Manipuri (mni) 8000 1000 1000 10000
Nepali (npi) 8000 1000 1000 10000
Odia (ory) 8000 1000 1000 10000
Punjabi (pan) 8000 1000 1000 10000
sanskrit(san) 8000 1000 1000 10000
Santali (sat) 8000 1000 1000 10000
Sindhi (snd) 8000 1000 1000 10000
Tamil (tam) 8000 1000 1000 10000
Telugu (tel) 8000 1000 1000 10000
Urdu (urd) 8000 1000 1000 10000
English (eng) 8000 1000 1000 10000
Total 184000 23000 23000 230000
Table 1: Data splits for ILID Benchmark Dataset
Figure 1: Workflow of the Proposed Language Identification System.
To ensure quality and consistency, the dataset undergoes several noise removal steps. The first step involves the elimination of duplicate, very short, and ungrammatical sentences. The next step employs an existing FastText Joulin et al. (2016b, a) based language identification model to remove sentences where the probability of detecting the language of the text is very low (we fix a threshold of 0.7 for this filtering process). This happens in code-mixed texts as many Indians are multilingual  [^2] https://en.wikipedia.org/wiki/List_of_languages_by_number_of_native_speakers_in_India and often use more than one language while writing.
The first approach is described in details.
2.1 Web Scraping
To build a strong language identification system for Indian languages, we carefully built a personalized set of 10,000 text samples for each language, covering 12 Indian languages. Details are available in Appendix A Table 4.
Text data is collected from a variety of publicly available and diverse sources, including Wikipedia dumps in respective languages for formal and structured text, news websites and blogs for professional and personal updates. All efforts are made to ensure the samples are multilingual in length, script (i.e., Devanagari, Bengali, Tamil, Odia etc.), domain, and style, reflecting the multilingual nature of India.
For a target website $W$, we define the scraping process as:
$S(W)=\bigcup_{p\in P_{W}}\phi(p)$ (1)
where $P_{W}$ is the set of target pages and $\phi$ is the extraction function that parses HTML content while preserving structural information. We implement adaptive throttling using:
$\Delta_{t}=\frac{1}{|W|}\sum_{i=1}^{|W|}\frac{\text{size}(w_{i})}{\text{bandwidth}}$ (2)
2.2 Data Cleaning Pipeline
In the data cleaning process we mainly work on removing extra spaces, extra special symbols, unicode normalization, breaking the paragraph into meaningful sentences, and after that we apply various tokenization methods. For sentence tokenization, we utilize the corresponding end sentence markers defined for respective languages. We use a regular expression based tokenizer  [^3] https://github.com/Pruthwik/Tokenizer_for_Indian_Languages for sentence and word tokenization. Our cleaning process transforms raw text $T$ to cleaned text $T^{\prime}$ through:
$T^{\prime}=\psi_{n}(\psi_{s}(T))$ (3)
where $\psi_{s}$ handles special characters and $\psi_{n}$ performs normalization. We use a sentence as a unit for the LID task. Each sentence is labeled with a language tag. ISO 639-2 codes  [^4] https://en.wikipedia.org/wiki/List_of_ISO_639-2_codes or 3 lettered language codes denote each language detailed in Appendix B Table 5. In the second approach, sentences are randomly sampled from the huge corpora Mujadia and Sharma (2024) that contain the information about the language. A similar exercise of appending the labels to the sentences as language identifiers and data splitting is performed on these samples.
3 ILID Model
The ILID model is a classifier that can categorize a piece of text into one of the 23 classes that includes English and 22 Indian languages. The architecture of the proposed system is shown in Figure 1. We implement three kinds of approaches for designing the classifiers.
3.1 Machine Learning Models
In this approach, each sentence is represented by a TF-IDF Sparck Jones (1972) vector in a bag-of-words setting. We utilize both word-level and character-level TF-IDF representations to enhance textual feature extraction. Word-level TF-IDF captures the importance of entire words across the corpus, which helps in preserving semantic meaning and understanding context at a higher level. However, word-level models may struggle with noisy or morphologically rich languages, spelling errors, or out-of-vocabulary terms. To address this, we also incorporate character-level TF-IDF, which analyzes smaller units such as character n-grams. This enables the model to recognize sub-word patterns and partial word structures, making it more resilient to variations and linguistic nuances. By combining both approaches, our system is able to capture both the global meaning (through words) and the finer-grained patterns (through characters), resulting in a more robust and language-agnostic representation, especially beneficial for multilingual or noisy text data. Eight different classifiers such as Support Vector Machine (SVM), Logistic Regression (LR), Random Forest (RF), Decision Tree (DT), Naive Bayes’, Stochastic Gradient Descent (SGD), K-Nearest Neighbors, and Adaboost are explored for this task. We develop different ensemble models consisting of 3, 4, and 5 classifiers and evaluate their performance. Due to paucity of space, we only report the five best performing ensemble models.
Model KNN DT RF SVM NB LogReg SGD F1 (Dev) F1 (Test)
Voting-1 0.99 0.96
Voting-2 0.99 0.96
Voting-3 0.99 0.96
Voting-4 0.99 0.96
Voting-5 0.99 0.99
MuRIL 0.96 0.96
FastText 0.96 0.96
Table 2: Comparison of Voting Classifiers, MuRIL, and FastText on Dev and Test sets
Model KNN DT RF SVM NB LogReg SGD ILID IndicLID
Voting-1 0.98 0.94
Voting-2 0.98 0.96
Voting-3 0.98 0.95
Voting-4 0.98 0.95
Voting-5 0.98 0.97
MuRIL 0.98 0.94
FastText 0.98 0.92
Table 3: Comparison of Voting Classifiers, MuRIL, and FastText on Other Benchmarks
3.2 FastText Classifier
FastText Joulin et al. (2016b, a) utilizes word embeddings Bojanowski et al. (2017) composed of character n-grams that can better represent rare words and orthographic similarities. FastText classifier is a linear classifier and is very fast at computation.
4 Pretrained BERT Model
Pretrained subword based contextual language models such as BERT Devlin et al. (2019), XLM Conneau et al. (2020), RoBERTa Zhuang et al. (2021) have proven to be very effective in text classification tasks or generally natural language understanding tasks. One variant of BERT, MuRIL Khanuja et al. (2021) is pretrained on Indian languages data. Hence, we fine-tune the MuRIL pretrained model on the ILID train set for our task.
5 Experimental Details
The machine learning models have been implemented using the Scikit-learn Pedregosa et al. (2011) framework. Similarly, the FastText library  [^5] https://fasttext.cc/docs/en/supervised-tutorial.html is utilized to implement the classification on the languages of the texts. The MuRIL models is fine-tuned using the Hugginggace transformer Wolf et al. (2019) framework. We will upload the models with MIT license once the paper gets accepted.
6 Evaluation Metrics
All models have been evaluated using the macro F1 scores. The macro F1 score averages the F1 scores across all the languages. In order to evaluate the score, the precision and recall scores are also computed at the macro level. Each ensemble employs majority voting mechanism to determine the predicted labels.
7 Results and Discussion
The ensemble machine learning models perform better than the individual models. The performance of the ensembles is also superior to the FastText and fine-tuned MuRIL models. The best five performing ensembles along with the FastText and fine-tuned MuRIL models are presented in Table 2 that represent the macro F1 score for each model. The results for the individual models are shown in Tables 6 and 7. Although the scores are encouraging, the performance drops in languages that share a script such as Devanagari for Hindi, Maithili, Marathi, Konkani, and Sanskrit, Arabic for Kashmiri, Urdu, and Sindhi. We compare our models with the state-of-the-art IndicLID model Madhani et al. (2023) on the Bhasha-Abhijnaanam Madhani et al. (2023) benchmark where our model perform comparatively. Our ILID model perform better on the languages where data are scraped while the performance dips for languages that are created from other external resources.
8 Conclusion
In this paper, we create the ILID benchmark dataset for English and Indian languages and develop baseline language identification models. The dataset and models will serve as foundation tools mostly for the resource poor languages that are part of this benchmark. We hope that this will act as an impetus for future NLP research in these languages.
Appendix A Languages Used in Scraping and Sampling
As explained in 2, based on their online availability, we used two approaches of scraping and sampling from an existing corpus for the dataset creation. The languages are shown in Table 4.
Scraping Sampling
Assamese Bodo
Bengali Dogri
Gujarati Kashmiri
Hindi Konkani
Kannada Maithili
Malayalam Manipuri
Marathi Nepali
Oriya Sanskrit
Punjabi Santhali
Tamil Sindhi
Telugu
Urdu
English
Table 4: Languages Used in Scraping and Sampling
Appendix B ISO 639-2 Codes
We utilize ISO 639-2 codes as the language labels for this task.
Language ISO 639-2 Code
Assamese asm
Bangla ben
Bodo brx
Dogri doi
Gujarati guj
Hindi hin
Kannada kan
Kashmiri kas
Konkani gom
Maithili mai
Malayalam mal
Manipuri mni
Marathi mar
Nepali npi
Oriya ory
Punjabi pan
Sanskrit san
Santali sat
Sindhi snd
Tamil tam
Telugu tel
Urdu urd
English eng
Table 5: Language Full Name to ISO 639-2 Language Codes Mapping
Appendix C Results for Individual Languages
The language wise performance of the models are shown in Tables 6 and 7.
Language LR DT RF SVM SGD k-NN NB ADABoost MURIL FastText
Assamese (asm) 0.99 0.97 0.99 0.99 0.98 0.91 0.98 0.77 0.99 0.99
Bengali (ben) 1.00 0.99 0.99 1.00 0.99 0.93 0.99 0.92 1.00 1.00
Bodo (brx) 0.96 0.92 0.96 0.97 0.94 0.90 0.96 0.68 0.97 0.96
Dogri (doi) 0.90 0.82 0.89 0.90 0.88 0.71 0.90 0.19 0.95 0.92
English (eng) 1.00 0.99 1.00 1.00 1.00 1.00 0.99 0.95 1.00 1.00
Konkani (gom) 0.97 0.85 0.95 0.97 0.93 0.80 0.96 0.29 0.98 0.98
Gujarati (guj) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
Hindi (hin) 0.99 0.93 0.99 0.99 0.98 0.83 0.94 0.79 1.00 0.99
Kannada (kan) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
Kashmiri (kas) 0.99 0.95 0.97 0.99 0.94 0.97 0.98 0.75 1.00 1.00
Maithili (mai) 0.87 0.78 0.84 0.86 0.84 0.67 0.84 0.26 0.92 0.90
Malayalam (mal) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
Marathi (mar) 0.98 0.84 0.96 0.98 0.93 0.76 0.94 0.52 0.99 0.99
Manipuri (mni) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 0.72 0.72
Nepali (npi) 0.92 0.84 0.90 0.91 0.86 0.76 0.89 0.28 0.94 0.92
Odia (ory) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 1.00 1.00
Punjabi (pan) 1.00 0.99 1.00 1.00 0.99 0.99 0.99 0.91 1.00 1.00
Sanskrit (san) 0.97 0.92 0.99 0.98 0.97 0.84 0.94 0.78 0.99 0.99
Santali (sat) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.95 0.71 0.71
Sindhi (snd) 0.94 0.86 0.92 0.93 0.90 0.80 0.94 0.24 0.96 0.94
Tamil (tam) 1.00 0.99 1.00 1.00 0.99 0.99 0.99 0.95 1.00 1.00
Telugu (tel) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.91 1.00 1.00
Urdu (urd) 0.99 0.94 0.96 0.99 0.96 0.97 0.98 0.76 1.00 1.00
Table 6: Dev dataset report: word+char combined Classifier performance (F1-scores) across 22 languages. Bold indicates top-performing model(s) per language.
Language LR DT RF SVM SGD k-NN NB AdaBoost MuRIL FastText
Assamese (asm) 0.99 0.96 0.98 0.99 0.98 0.91 0.97 0.78 0.99 0.99
Bengali (ben) 1.00 0.98 0.99 1.00 0.99 0.93 0.99 0.93 1.00 1.00
Bodo (brx) 0.95 0.90 0.96 0.96 0.95 0.89 0.96 0.69 0.97 0.96
Dogri (doi) 0.91 0.82 0.91 0.90 0.88 0.71 0.91 0.21 0.94 0.92
English (eng) 1.00 1.00 0.99 1.00 0.99 1.00 0.99 0.96 1.00 1.00
Konkani (gom) 0.97 0.86 0.94 0.97 0.92 0.81 0.95 0.31 0.98 0.97
Gujarati (guj) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
Hindi (hin) 0.99 0.93 0.98 0.99 0.97 0.84 0.94 0.74 0.99 0.99
Kannada (kan) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
Kashmiri (kas) 0.98 0.93 0.97 0.98 0.94 0.98 0.98 0.74 1.00 0.99
Maithili (mai) 0.85 0.78 0.84 0.85 0.81 0.67 0.82 0.26 0.91 0.89
Malayalam (mal) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
Marathi (mar) 0.97 0.85 0.95 0.98 0.92 0.79 0.94 0.50 0.98 0.98
Manipuri (mni) 1.00 1.00 1.00 1.00 1.00 0.99 1.00 0.89 0.73 0.72
Nepali (npi) 0.90 0.84 0.90 0.90 0.86 0.76 0.89 0.28 0.94 0.91
Odia (ory) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 1.00 1.00
Punjabi (pan) 1.00 0.99 0.99 1.00 0.99 0.99 1.00 0.91 1.00 1.00
Sanskrit (san) 0.98 0.92 0.99 0.98 0.97 0.84 0.94 0.79 0.98 0.98
Santali (sat) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.94 0.72 0.71
Sindhi (snd) 0.94 0.87 0.93 0.93 0.90 0.80 0.92 0.23 0.96 0.94
Tamil (tam) 1.00 0.99 1.00 1.00 0.99 1.00 1.00 0.96 1.00 1.00
Telugu (tel) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.92 1.00 1.00
Urdu (urd) 0.98 0.92 0.97 0.99 0.96 0.98 0.98 0.75 1.00 0.99
Table 7: Test dataset report showing word+char combined classifier performance (F1-scores) across 22 languages. Bold indicates top-performing model(s) per language.

### v2

@@
 Abstract
-Language Identification task is a crucial fundamental step in NLP.
+The language identification task is a crucial fundamental step in NLP.
 Often it serves as a pre-processing step for widely used NLP applications such as multilingual machine translation, information retrieval, question and answering, and text summarization.
@@
 This becomes even harder in case of diverse Indian languages that exhibit lexical and phonetic similarities, but have distinct differences.
-Many Indian languages share the same script making the task even more challenging.
-In this paper, we release a dataset of 230K sentences consisting of English and all 22 official Indian languages labeled with their language identifiers where data in most languages are newly created.
-We also develop and release robust baseline models using state-of-the-art approaches in machine learning and deep learning that can aid the research in this field.
-Our baseline models are comparable to the state-of-the-art models for the language identification task.
+Many Indian languages share the same script, making the task even more challenging.
+Taking all these challenges into account, we develop and release a dataset of 250K sentences consisting of 23 languages including English and all 22 official Indian languages labeled with their language identifiers, where data in most languages are newly created.
+We also develop and release baseline models using state-of-the-art approaches in machine learning and fine-tuning pre-trained transformer models.
+Our models outperforms the state-of-the-art pre-trained transformer models for the language identification task.
+The dataset and the codes are available at https://yashingle-ai.github.io/ILID/ and in Huggingface open source libraries.
 1 Introduction
@@
 With the increase in mobile and internet users in India, the need for providing services in native Indian languages is more pressing than ever.
-Hence, it becomes essential to create robust LID tools to cater to the Indian users.
-As the volume and variety of data keeps growing, we need to include data from different domains while creating sound benchmarks.
-Keeping this motivation in mind, we set out to create an Indian Language IDentification (ILID) benchmark including English and 22 official languages consisting of a total 230K sentences labeled with their language markers.
-The main contribution of our paper is two folds:
-  • We create ILID dataset consisting of newly created datasets in 13 languages and curated datasets in rest of the languages.
-  • We develop ILID baseline models exploring different machine learning and deep learning techniques.
+Hence, it becomes essential to create robust LID tools to cater to Indian users.
+As the volume and variety of data continue to grow, we need to include data from different domains while creating sound benchmarks.
+Keeping this motivation in mind, we set out to create an Indian Language Identification (ILID) benchmark including English and 22 official languages spanning 25 scripts consisting of a total of 250K sentences labeled with their language markers.
+The main contribution of our paper is two fold.
+  • We create the ILID dataset consisting of newly created datasets in 13 languages and curated datasets in the remaining ten languages.
+  • We develop ILID baseline models using different machine learning and deep learning techniques.
+  • We perform a comparative analysis of our models with the state-of-the-art model.
+2 Related Work
+Language Identification (LID) is a fundamental task in Natural Language Processing (NLP), aiming to determine the language of a given text or speech segment.
+Early approaches to LID primarily relied on statistical methods, most notably character $n$-gram models Cavnar and Trenkle (1994).
+These methods leverage the unique frequency distributions of character sequences within different languages to classify unknown texts.
+Language Identification becomes more crucial in the domain of Cross Lingual Information Retrieval where the query is written in native scripts Bosca and Dini (2010).
+Similarly, spoken language identificationDehak et al.
+(2011); Gonzalez-Dominguez et al.
+(2014) is a very essential task in multilingual speech processing that can further impact in developing applications such as machine translation and speech recognition.
+In this paper, we are dealing with the language identification task in written texts.
+With machine learning approaches such as Support Vector Machines (SVMs) Dunning (1994), Naive Bayes classifiers Elworthy (1998), decision trees, random forests, and gradient boosting gaining widespread usage, LID systems were also modeled to leverage these techniques with the same $n$-gram features and other lexical or language specific hand crafted features.
+While effective for well-resourced languages with distinct orthographies, the performance of the statistical systems can degrade when dealing with short texts, noisy data, or languages sharing common scripts.
+Deep learning methods such as Convolutional Neural Networks (CNNs) Kim (2014) and Recurrent Neural Networks (RNNs) Hochreiter and Schmidhuber (1997) (particularly LSTMs and GRUs) are often trained on distributionally similar word or character n-gram or single character embeddings vectors to better capture intricate language complexity and handle out-of-vocabulary words more robustly Conneau et al.
+(2017).
+The current approaches include fine-tuning pre-trained transformer Vaswani et al.
+(2017) models using subwords Sennrich et al.
+(2016); Kudo and Richardson (2018); Song et al.
+(2021).
+Many of the language identification tasks are modeled as token classification tasks that are helpful in code-mixed settings.
+The task of language identification for Indian languages is even more challenging as they belong to different and diverse language families (Indo-Aryan, Dravidian, Tibeto-Burman, Austroasiatic), and written in over a dozen distinct scripts.
+Multiple Indian languages often share the same script.
+For example; Hindi, Marathi, Nepali, and Sanskrit are all primarily written in the Devanagari script.
+This script overlap requires models to distinguish languages based on lexical, morphological, or syntactic features rather than solely on orthographic cues.
+Another prevalent issue in Indian language contexts is code-mixing and code-switching, where speakers or writers regularly use words or phrases from multiple languages within a single utterance or text Gambäck and Das (2017).
+This phenomenon is common in informal communication and social media Barman et al.
+(2014), making it difficult for traditional LID systems to accurately identify the primary language or even segment code-mixed segments.
+Research in this area often focuses on fine-grained LID, aiming to identify language at the word or phrase level Bhat et al.
+(2014), rather than just document or sentence level identification.
+The language identification task for Indian languages also followed the similar line of research using rule-based and machine learning based approaches.
+More advanced approaches fine-tune pre-train transformers trained on Indian languages Kumar et al.
+(2023); Agarwal et al.
+(2023); Madhani et al.
+(2023a).
+Madhani et al.
+(2023a) also transliterate Madhani et al.
+(2023b) the native scripts into roman scripts for uniformity across languages and better efficiency in the LID systems.
+Despite these advancements, robust and fine-grained LID for all Indian languages, especially in code-mixed scenarios and for low-resource languages, remains an active area of research.
+In this paper, we limit the scope of identifying the language at a sentence level, not at the constituent token level.
 3 ILID Dataset Creation
 Indian Language Identification (ILID) dataset is created using two approaches.
-We include English and all 22 official Indian languages  [^1] https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India widely used in India.
+We include English and the 22 official Indian languages  [^1] https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India [^2] https://en.wikipedia.org/wiki/Linguistic_Survey_of_India widely used in India.
 The first approach utilizes web scraping for the languages in which the digital presence is significant (details are given in the Appendix).
-For each of these languages, we collect 10,000 different sentences from diverse sources such as government websites, newspapers, books, and other public materials ensuring varying degrees of linguistic complexity.
-The second approach samples sentences from an existing massive monolingual and parallel corpora for Indian languages Mujadia and Sharma (2024).
+For each of these languages, we collect 10,000 different sentences from various sources such as government websites, newspapers, books, and other public materials, ensuring varying degrees of linguistic complexity.
+The second approach samples sentences from Bhashaverse Mujadia and Sharma (2024), an existing massive monolingual and parallel corpora for Indian languages.
 The details of the dataset are presented in Table 1.
 The data in each language is split into 80:10:10 ratio to create train, dev, and test sets.
+The dataset is available at https://huggingface.co/datasets/yash-ingle/ILID_Indian_Language_Identification_Dataset.
 Language #Train #Dev #Test #Total
@@
 Marathi (mar) 8000 1000 1000 10000
-Manipuri (mni) 8000 1000 1000 10000
+Manipuri Bengali Script (mni_Beng) 8000 1000 1000 10000
+Manipuri Meitei Script (mni_Mtei) 8000 1000 1000 10000
 Nepali (npi) 8000 1000 1000 10000
@@
 Santali (sat) 8000 1000 1000 10000
-Sindhi (snd) 8000 1000 1000 10000
+Sindhi Devnagari(snd_Deva) 8000 1000 1000 10000
+Sindhi Perso-Arabic(snd_Arab) 8000 1000 1000 10000
 Tamil (tam) 8000 1000 1000 10000
@@
 English (eng) 8000 1000 1000 10000
-Total 184000 23000 23000 230000
+Total 200000 25000 25000 250000
 Table 1: Data splits for ILID Benchmark Dataset
@@
 (2016b, a) based language identification model to remove sentences where the probability of detecting the language of the text is very low (we fix a threshold of 0.7 for this filtering process).
-This happens in code-mixed texts as many Indians are multilingual  [^2] https://en.wikipedia.org/wiki/List_of_languages_by_number_of_native_speakers_in_India and often use more than one language while writing.
-The first approach is described in details.
+This happens in code-mixed texts as many Indians are multilingual and often use more than one language while writing.
+The approaches are described in detail.
 3.1 Web Scraping
-To build a strong language identification system for Indian languages, we carefully built a personalized set of 10,000 text samples for each language, covering 12 Indian languages.
-Details are available in Appendix A Table 4.
+To build a strong language identification system for Indian languages, we carefully built a personalized set of 10,000 text samples for each language, covering 13 Indian languages.
 Text data is collected from a variety of publicly available and diverse sources, including Wikipedia dumps in respective languages for formal and structured text, news websites and blogs for professional and personal updates.
@@
 3.2 Data Cleaning Pipeline
-In the data cleaning process we mainly work on removing extra spaces, extra special symbols, unicode normalization, breaking the paragraph into meaningful sentences, and after that we apply various tokenization methods.
+In the data cleaning process, we mainly work on removing extra spaces, extra special symbols, unicode normalization, breaking the paragraph into meaningful sentences, and after that we apply various tokenization methods.
 For sentence tokenization, we utilize the corresponding end sentence markers defined for respective languages.
-We use a regular expression based tokenizer  [^3] https://github.com/Pruthwik/Tokenizer_for_Indian_Languages for sentence and word tokenization.
+We use a regular expression based tokenizer  [^3] https://github.com/Pruthwik/Tokenizer_for_Indian_Languages for sentence and word tokenization that is specifically designed to handle texts in Indian languages.
 Our cleaning process transforms raw text $T$ to cleaned text $T^{\prime}$ through:
@@
 Each sentence is labeled with a language tag.
-ISO 639-2 codes  [^4] https://en.wikipedia.org/wiki/List_of_ISO_639-2_codes or 3 lettered language codes denote each language detailed in Appendix B Table 5.
-In the second approach, sentences are randomly sampled from the huge corpora Mujadia and Sharma (2024) that contain the information about the language.
+ISO 639-2 codes  [^4] https://en.wikipedia.org/wiki/List_of_ISO_639-2_codes or 3 lettered language codes denote each language detailed in Table 3.
+3.3 Sampling From Existing Corpora
+In the second approach, sentences are randomly sampled from Bhashaverse Mujadia and Sharma (2024) that contain a huge collection of monolingual corpora of different Indian languages.
+In this corpora, Manipuri, and Sindhi have sentences in two different scripts each.
+Manipuri is available in Bengali and Meitei scripts, whereas Sindhi is written in Devanagari and Perso-Arabic scripts.
+We have sampled sentences using all available scripts from these two languages.
+This increases the total number of languages with different scripts to 25 also presented in Table 4.
 A similar exercise of appending the labels to the sentences as language identifiers and data splitting is performed on these samples.
-3 ILID Model
-The ILID model is a classifier that can categorize a piece of text into one of the 23 classes that includes English and 22 Indian languages.
-The architecture of the proposed system is shown in Figure 1.
-We implement three kinds of approaches for designing the classifiers.
-3.1 Machine Learning Models
-In this approach, each sentence is represented by a TF-IDF Sparck Jones (1972) vector in a bag-of-words setting.
-We utilize both word-level and character-level TF-IDF representations to enhance textual feature extraction.
-Word-level TF-IDF captures the importance of entire words across the corpus, which helps in preserving semantic meaning and understanding context at a higher level.
-However, word-level models may struggle with noisy or morphologically rich languages, spelling errors, or out-of-vocabulary terms.
-To address this, we also incorporate character-level TF-IDF, which analyzes smaller units such as character n-grams.
-This enables the model to recognize sub-word patterns and partial word structures, making it more resilient to variations and linguistic nuances.
-By combining both approaches, our system is able to capture both the global meaning (through words) and the finer-grained patterns (through characters), resulting in a more robust and language-agnostic representation, especially beneficial for multilingual or noisy text data.
-Eight different classifiers such as Support Vector Machine (SVM), Logistic Regression (LR), Random Forest (RF), Decision Tree (DT), Naive Bayes’, Stochastic Gradient Descent (SGD), K-Nearest Neighbors, and Adaboost are explored for this task.
-We develop different ensemble models consisting of 3, 4, and 5 classifiers and evaluate their performance.
-Due to paucity of space, we only report the five best performing ensemble models.
-Model KNN DT RF SVM NB LogReg SGD F1 (Dev) F1 (Test)
-Voting-1 0.99 0.96
-Voting-2 0.99 0.96
-Voting-3 0.99 0.96
-Voting-4 0.99 0.96
-Voting-5 0.99 0.99
-MuRIL 0.96 0.96
-FastText 0.96 0.96
-Table 2: Comparison of Voting Classifiers, MuRIL, and FastText on Dev and Test sets
-Model KNN DT RF SVM NB LogReg SGD ILID IndicLID
-Voting-1 0.98 0.94
-Voting-2 0.98 0.96
-Voting-3 0.98 0.95
-Voting-4 0.98 0.95
-Voting-5 0.98 0.97
-MuRIL 0.98 0.94
-FastText 0.98 0.92
-Table 3: Comparison of Voting Classifiers, MuRIL, and FastText on Other Benchmarks
-3.2 FastText Classifier
-FastText Joulin et al.
-(2016b, a) utilizes word embeddings Bojanowski et al.
-(2017) composed of character n-grams that can better represent rare words and orthographic similarities.
-FastText classifier is a linear classifier and is very fast at computation.
-4 Pretrained BERT Model
-Pretrained subword based contextual language models such as BERT Devlin et al.
-(2019), XLM Conneau et al.
-(2020), RoBERTa Zhuang et al.
-(2021) have proven to be very effective in text classification tasks or generally natural language understanding tasks.
-One variant of BERT, MuRIL Khanuja et al.
-(2021) is pretrained on Indian languages data.
-Hence, we fine-tune the MuRIL pretrained model on the ILID train set for our task.
-5 Experimental Details
-The machine learning models have been implemented using the Scikit-learn Pedregosa et al.
-(2011) framework.
-Similarly, the FastText library  [^5] https://fasttext.cc/docs/en/supervised-tutorial.html is utilized to implement the classification on the languages of the texts.
-The MuRIL models is fine-tuned using the Hugginggace transformer Wolf et al.
-(2019) framework.
-We will upload the models with MIT license once the paper gets accepted.
-6 Evaluation Metrics
-All models have been evaluated using the macro F1 scores.
-The macro F1 score averages the F1 scores across all the languages.
-In order to evaluate the score, the precision and recall scores are also computed at the macro level.
-Each ensemble employs majority voting mechanism to determine the predicted labels.
-7 Results and Discussion
-The ensemble machine learning models perform better than the individual models.
-The performance of the ensembles is also superior to the FastText and fine-tuned MuRIL models.
-The best five performing ensembles along with the FastText and fine-tuned MuRIL models are presented in Table 2 that represent the macro F1 score for each model.
-The results for the individual models are shown in Tables 6 and 7.
-Although the scores are encouraging, the performance drops in languages that share a script such as Devanagari for Hindi, Maithili, Marathi, Konkani, and Sanskrit, Arabic for Kashmiri, Urdu, and Sindhi.
-We compare our models with the state-of-the-art IndicLID model Madhani et al.
-(2023) on the Bhasha-Abhijnaanam Madhani et al.
-(2023) benchmark where our model perform comparatively.
-Our ILID model perform better on the languages where data are scraped while the performance dips for languages that are created from other external resources.
-8 Conclusion
-In this paper, we create the ILID benchmark dataset for English and Indian languages and develop baseline language identification models.
-The dataset and models will serve as foundation tools mostly for the resource poor languages that are part of this benchmark.
-We hope that this will act as an impetus for future NLP research in these languages.
-Appendix A Languages Used in Scraping and Sampling
-As explained in 2, based on their online availability, we used two approaches of scraping and sampling from an existing corpus for the dataset creation. The languages are shown in Table 4.
+The language specifics using both approaches are shown in Table 2.
 Scraping Sampling
@@
 Table 2: Languages Used in Scraping and Sampling
-Appendix B ISO 639-2 Codes
-We utilize ISO 639-2 codes as the language labels for this task.
 Language ISO 639-2 Code
@@
 Malayalam mal
-Manipuri mni
+Manipuri Bengali Script mni_Beng
+Manipuri Meitei Scrip mni_Mtei
 Marathi mar
@@
 Santali sat
-Sindhi snd
+Sindhi Perso-Arabic snd_Arab
+Sindhi Devanagari snd_Deva
 Tamil tam
@@
 Table 3: Language Full Name to ISO 639-2 Language Codes Mapping
-Appendix C Results for Individual Languages
-The language wise performance of the models are shown in Tables 6 and 7.
-Language LR DT RF SVM SGD k-NN NB ADABoost MURIL FastText
-Assamese (asm) 0.99 0.97 0.99 0.99 0.98 0.91 0.98 0.77 0.99 0.99
-Bengali (ben) 1.00 0.99 0.99 1.00 0.99 0.93 0.99 0.92 1.00 1.00
-Bodo (brx) 0.96 0.92 0.96 0.97 0.94 0.90 0.96 0.68 0.97 0.96
-Dogri (doi) 0.90 0.82 0.89 0.90 0.88 0.71 0.90 0.19 0.95 0.92
-English (eng) 1.00 0.99 1.00 1.00 1.00 1.00 0.99 0.95 1.00 1.00
-Konkani (gom) 0.97 0.85 0.95 0.97 0.93 0.80 0.96 0.29 0.98 0.98
-Gujarati (guj) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
-Hindi (hin) 0.99 0.93 0.99 0.99 0.98 0.83 0.94 0.79 1.00 0.99
-Kannada (kan) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
-Kashmiri (kas) 0.99 0.95 0.97 0.99 0.94 0.97 0.98 0.75 1.00 1.00
-Maithili (mai) 0.87 0.78 0.84 0.86 0.84 0.67 0.84 0.26 0.92 0.90
-Malayalam (mal) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
-Marathi (mar) 0.98 0.84 0.96 0.98 0.93 0.76 0.94 0.52 0.99 0.99
-Manipuri (mni) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 0.72 0.72
-Nepali (npi) 0.92 0.84 0.90 0.91 0.86 0.76 0.89 0.28 0.94 0.92
-Odia (ory) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 1.00 1.00
-Punjabi (pan) 1.00 0.99 1.00 1.00 0.99 0.99 0.99 0.91 1.00 1.00
-Sanskrit (san) 0.97 0.92 0.99 0.98 0.97 0.84 0.94 0.78 0.99 0.99
-Santali (sat) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.95 0.71 0.71
-Sindhi (snd) 0.94 0.86 0.92 0.93 0.90 0.80 0.94 0.24 0.96 0.94
-Tamil (tam) 1.00 0.99 1.00 1.00 0.99 0.99 0.99 0.95 1.00 1.00
-Telugu (tel) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.91 1.00 1.00
-Urdu (urd) 0.99 0.94 0.96 0.99 0.96 0.97 0.98 0.76 1.00 1.00
-Table 6: Dev dataset report: word+char combined Classifier performance (F1-scores) across 22 languages. Bold indicates top-performing model(s) per language.
-Language LR DT RF SVM SGD k-NN NB AdaBoost MuRIL FastText
-Assamese (asm) 0.99 0.96 0.98 0.99 0.98 0.91 0.97 0.78 0.99 0.99
-Bengali (ben) 1.00 0.98 0.99 1.00 0.99 0.93 0.99 0.93 1.00 1.00
-Bodo (brx) 0.95 0.90 0.96 0.96 0.95 0.89 0.96 0.69 0.97 0.96
-Dogri (doi) 0.91 0.82 0.91 0.90 0.88 0.71 0.91 0.21 0.94 0.92
-English (eng) 1.00 1.00 0.99 1.00 0.99 1.00 0.99 0.96 1.00 1.00
-Konkani (gom) 0.97 0.86 0.94 0.97 0.92 0.81 0.95 0.31 0.98 0.97
-Gujarati (guj) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
-Hindi (hin) 0.99 0.93 0.98 0.99 0.97 0.84 0.94 0.74 0.99 0.99
-Kannada (kan) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
-Kashmiri (kas) 0.98 0.93 0.97 0.98 0.94 0.98 0.98 0.74 1.00 0.99
-Maithili (mai) 0.85 0.78 0.84 0.85 0.81 0.67 0.82 0.26 0.91 0.89
-Malayalam (mal) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
-Marathi (mar) 0.97 0.85 0.95 0.98 0.92 0.79 0.94 0.50 0.98 0.98
-Manipuri (mni) 1.00 1.00 1.00 1.00 1.00 0.99 1.00 0.89 0.73 0.72
-Nepali (npi) 0.90 0.84 0.90 0.90 0.86 0.76 0.89 0.28 0.94 0.91
-Odia (ory) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 1.00 1.00
-Punjabi (pan) 1.00 0.99 0.99 1.00 0.99 0.99 1.00 0.91 1.00 1.00
-Sanskrit (san) 0.98 0.92 0.99 0.98 0.97 0.84 0.94 0.79 0.98 0.98
-Santali (sat) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.94 0.72 0.71
-Sindhi (snd) 0.94 0.87 0.93 0.93 0.90 0.80 0.92 0.23 0.96 0.94
-Tamil (tam) 1.00 0.99 1.00 1.00 0.99 1.00 1.00 0.96 1.00 1.00
-Telugu (tel) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.92 1.00 1.00
-Urdu (urd) 0.98 0.92 0.97 0.99 0.96 0.98 0.98 0.75 1.00 0.99
-Table 7: Test dataset report showing word+char combined classifier performance (F1-scores) across 22 languages. Bold indicates top-performing model(s) per language.
+lang #sents #words #chars avg_word_len avg_sent_len #unique_words TTR_words
+mai 10000 65884 288223 3.526 28.822 6743 0.102
+eng 10000 225036 1052283 3.721 105.228 22709 0.101
+kan 10000 122400 861095 6.117 86.109 34644 0.283
+san 10000 130353 858829 5.665 85.883 50506 0.387
+gom 10000 58282 296807 4.264 29.681 8104 0.139
+hin 10000 106402 476919 3.576 47.692 8451 0.079
+mar 10000 89232 526021 5.007 52.602 19470 0.218
+asm 10000 64251 311313 4.001 31.131 7196 0.112
+urd 10000 97233 387031 3.083 38.703 19987 0.206
+tel 10000 81281 524216 5.572 52.422 19554 0.241
+tam 10000 144307 1018268 6.126 101.827 45912 0.318
+snd_Arab 10000 113249 479446 3.322 47.945 7409 0.065
+ben 10000 170048 877189 4.217 87.719 30007 0.176
+snd_Deva 10000 70743 312570 3.56 31.257 6249 0.088
+ory 10000 91919 526927 4.841 52.693 5400 0.059
+npi 10000 59095 296473 4.186 29.647 7966 0.135
+mni_Mtei 10000 73043 318189 3.493 31.819 7245 0.099
+mal 10000 117447 985898 7.48 98.59 37010 0.315
+doi 10000 75469 313613 3.288 31.361 5904 0.078
+pan 10000 168450 721285 3.341 72.129 18032 0.107
+brx 10000 65717 343731 4.383 34.373 7828 0.119
+sat 10000 78554 360814 3.72 36.081 4067 0.052
+mni_Beng 10000 186109 1208020 5.545 120.802 27350 0.147
+guj 10000 175966 911173 4.235 91.117 28195 0.16
+kas 10000 69270 297326 3.437 29.733 7752 0.112
+Table 4: Language Wise Corpus Statistics Where the column TTR_words refers to the Type Token Ratio for Words
+Model KNN DT RF SVM NB LogReg SGD ILID Bh-Ab
+Dev Test
+Voting-1 1.00 0.96 1.00
+Voting-2 1.00 0.96 1.00
+Voting-3 1.00 0.96 1.00
+Voting-4 0.99 0.96 1.00
+Voting-5 0.99 0.99 1.00
+MuRIL 0.96 0.96 0.9
+FastText 0.96 0.96 0.9
+Table 5: Comparison of Voting Classifiers, MuRIL, and FastText on ILID dataset and Bhasha-Abhijnaanam (Bh-Ab) dataset. Highest scores are marked in bold.
+3.4 Corpora Statistics
+After collecting and cleaning the dataset, we compute several corpora statistics.
+These statistics are excellent indicators of the levels and complexities of each language.
+The properties of the corpora considered include the total number of words, the total number of characters, the total number of unique words, average sentence length, the average word length, and the type / token ratio textinspector (2025) which are markers of the representativeness of the corpora.
+These statistics are shown in Table 4.
+4 ILID Model
+The ILID model is a classifier that can categorize a piece of text into one of the 25 classes that includes English and 22 Indian languages.
+The architecture of the proposed system is shown in Figure 1.
+We implement three types of approach for designing the classifiers.
+4.1 Machine Learning Models
+In this approach, each sentence is represented by a TF-IDF Sparck Jones (1972) vector in a bag-of-words setting.
+We utilize both word-level and character-level TF-IDF representations to enhance textual feature extraction.
+4.1.1 Word-Level TF-IDF Based Classification
+[…]
+4.1.2 Character-Level TF-IDF Based Classification
+To overcome the limitations of word-level models, we introduce character-level TF-IDF representations using character n-grams ranging from 2 to 6.
+This approach captures sub-word structures, prefixes, suffixes, and root forms that are particularly effective in handling typographical errors, morphological variations, and out-of-vocabulary words.
+Character-level features are especially helpful in language identification tasks involving noisy or informal text, enabling the model to learn finer linguistic patterns.
+4.1.3 Combined Feature-Based Classification
+To leverage the strengths of both word and character level features, we develop a combined TF-IDF representation by concatenating the two feature sets into a single vector.
+This hybrid approach provides a more robust and language-agnostic representation.
+We train eight traditional machine learning classifiers—Support Vector Machine (SVM), Logistic Regression (LR), Random Forest (RF), Decision Tree (DT), Naive Bayes (NB), Stochastic Gradient Descent (SGD), and K-Nearest Neighbors (KNN)—on these features.
+Furthermore, we explore various ensemble models involving combinations of 3, 4, and 5 diverse base classifiers from the pool of classifiers.
+Due to space constraints, only the top five performing ensembles are selected based on F1 scores on the development and test sets, are presented in Table 5.
+4.2 FastText Classifier
+FastText Joulin et al.
+(2016b, a) utilizes word embeddings Bojanowski et al.
+(2017) composed of character n-grams that better represent rare words and orthographic similarities.
+FastText classifier is a linear classifier and is very fast in computation.
+It can generate embeddings for out-of-vocabulary (OOV) words, which are common in user-generated or noisy text data.
+4.3 Pretrained BERT Model
+Pretrained subword based contextual language models such as BERT Devlin et al.
+(2019), XLM Conneau et al.
+(2020), RoBERTa Zhuang et al.
+(2021) have proven to be very effective in text classification tasks or generally natural language understanding tasks.
+One variant of BERT, MuRIL Khanuja et al.
+(2021) is pretrained on large amounts of corpora in Indian languages.
+The pre-training data is also augmented with translated and transliterated texts that makes the model capture cross-lingual and code-mixed embeddings more efficiently.
+Hence, we fine-tune the MuRIL pretrained model on the ILID train set for our task.
+5 Experimental Details
+The machine learning models have been implemented using the Scikit-learn Pedregosa et al.
+(2011) framework.
+Similarly, the FastText library  [^5] https://fasttext.cc/docs/en/supervised-tutorial.html is utilized to implement the language identifiers of the texts modeled as text classifiers Joulin et al.
+(2016b, a).
+The MuRIL Khanuja et al.
+(2021) models are fine-tuned using the Huggingface transformer Wolf et al.
+(2019) framework on an NVIDIA H100 GPU with 94GB RAM.
+The batch size, the number of epochs, the learning rate, maximum sequence length, and weight decay are set to 32, 10, 0.00002, 256, and 0.01 respectively.
+The fine-tuned MuRIL model is available at https://huggingface.co/pruthwik/ilid-muril-model.
+6 Evaluation Metrics
+All models have been evaluated using the macro F1 scores.
+The macro F1 score averages the F-1 scores across all languages.
+In order to evaluate the score the F1-score of each language, the precision and recall scores are also computed at the macro level.
+Each ensemble employs a voting mechanism to determine the predicted labels.
+Two types of voting mechanism are used: soft and hard.
+Hard voting is based on majority voting among the classifiers present in the ensemble.
+This type of voting is performed when the SVM classifier is part of the ensemble because SVM is inherently a non-probabilistic classifier.
+Soft voting aggregates the probability estimates of individual classifiers in the ensemble and the class with argmax of the sums of the predicted probabilities is chosen as the label.
+Ensembles where SVM is not included are evaluated using soft voting.
+7 Results and Discussion
+The ensemble machine learning models perform better than the individual models.
+The performance of the ensembles is also superior to the FastText and fine-tuned MuRIL models.
+The best five performing ensembles along with the FastText and fine-tuned MuRIL models are presented in Table 5 that represent the macro F1 score for each model.
+Although the scores are encouraging, the performance drops in languages that share a script such as Devanagari for Hindi, Maithili, Marathi, Konkani, and Sanskrit, Arabic for Kashmiri, Urdu, and Sindhi.
+We compare our models with the state-of-the-art IndicLID model Madhani et al.
+(2023a) on the Bhasha-Abhijnaanam Madhani et al.
+(2023a) benchmark of 88K sentences where our model outperforms the IndicLID model Madhani et al.
+(2023a).
+The F1-scores of the IndicLID model on the Bhasha-Abhijnaanam dataset is 0.98 as reported in the paper.
+We could not run their fine-tuned IndicBERT model while the IndicLID FTN model has a macro F1-score of 0.88.
+Our deep learning baselines perform better on the languages where data are scraped, while the performance dips for languages that are created from other external resources.
+7.1 Results for Individual Languages
+The language-wise performance of the models is shown in Tables 6.
+We can observe from the table that the performance drops for extremely low resource languages such as Bodo, Dogri, Maithili, and Sindhi.
+Deep learning models suffer in identifying those languages for which they have scarcity during pre-training.
+MuRIL Khanuja et al.
+(2021) outperforms every other model in languages on which the model was pre-trained.
+8 Conclusion
+In this paper, we present the ILID benchmark, a manually curated dataset for 23 languages, which includes English and 22 Indian languages containing 250K sentences.
+It is specifically designed for Indian language identification.
+Along with the dataset, we develop several machine learning an deep learning models that perform well, even with limited training data.
+This makes them especially suitable for low-resource languages.
+To check the consistency and reliability of our ML models, we compare them with deep learning models like MuRIL and FastText.
+The ILID dataset and models together provide important resources for advancing multilingual NLP research in Indian languages.
+We hope this work encourages more exploration and development in this area, which has been overlooked.
+Lang LR DT RF SVM SGD KNN NB Muril Fasttext
+Dev Test Dev Test Dev Test Dev Test Dev Test Dev Test Dev Test Dev Test Dev Test
+asm 0.98 0.99 0.96 0.96 0.98 0.98 0.98 0.99 0.97 0.98 0.90 0.96 0.97 0.98 0.99 0.99 0.96 0.96
+ben 1.00 1.00 0.97 0.98 0.99 0.99 1.00 1.00 0.99 0.99 0.91 0.97 0.99 0.99 1.00 1.00 0.98 0.98
+brx 0.96 0.95 0.91 0.89 0.96 0.94 0.96 0.95 0.94 0.95 0.85 0.93 0.96 0.95 0.91 0.91 0.92 0.91
+doi 0.92 0.92 0.83 0.82 0.90 0.91 0.91 0.92 0.88 0.88 0.68 0.81 0.90 0.91 0.93 0.92 0.91 0.91
+eng 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00 0.99 1.00 1.00 0.99 1.00 1.00 0.99 0.99
+gom 0.97 0.97 0.84 0.83 0.94 0.95 0.96 0.97 0.90 0.91 0.78 0.90 0.94 0.96 0.97 0.97 0.91 0.92
+guj 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.95 1.00 1.00 1.00 0.99 0.99
+hin 0.99 0.98 0.93 0.92 0.98 0.97 0.99 0.99 0.97 0.97 0.80 0.92 1.00 0.95 0.99 0.99 0.98 0.98
+kan 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.98 1.00 1.00 1.00 0.98 0.97
+kas 0.99 0.98 0.94 0.94 0.97 0.98 0.99 0.99 0.94 0.95 0.96 0.96 0.85 0.98 0.99 1.00 0.97 0.96
+mai 0.90 0.86 0.83 0.78 0.84 0.84 1.00 0.85 0.83 0.82 0.65 0.73 0.84 0.85 0.89 0.88 0.87 0.86
+mal 1.00 1.00 0.99 1.00 1.00 1.00 0.86 1.00 1.00 1.00 1.00 1.00 0.95 1.00 1.00 1.00 0.94 0.94
+mar 0.99 0.97 1.00 0.84 0.96 0.96 1.00 0.98 0.91 0.91 0.76 0.91 1.00 0.95 0.98 0.98 0.94 0.94
+mni_Beng 1.00 1.00 0.99 0.99 1.00 1.00 0.99 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99
+mni_Mtei 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.91 1.00 0.69 0.71 0.99 0.98
+npi 0.92 0.91 0.85 0.84 0.91 0.90 1.00 0.91 0.87 0.86 0.75 0.83 1.00 0.91 0.93 0.92 0.88 0.87
+ory 1.00 1.00 1.00 1.00 1.00 1.00 0.92 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
+pan 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00 0.99 1.00 0.97 1.00 1.00 1.00 0.96 0.97
+san 1.00 1.00 0.95 0.95 1.00 1.00 1.00 1.00 0.96 0.97 0.60 0.96 1.00 0.97 1.00 1.00 0.98 0.98
+sat 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99 0.92 1.00 0.70 0.72 0.99 0.99
+snd_arab 1.00 1.00 0.97 0.98 0.99 1.00 1.00 1.00 0.99 1.00 0.99 0.97 0.99 0.98 1.00 1.00 1.00 0.99
+snd_deva 0.94 0.93 0.88 0.85 0.92 0.91 0.94 0.92 0.89 0.90 0.80 0.88 0.92 0.92 0.94 0.94 0.92 0.92
+tam 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 0.99 0.99 0.90 1.00 0.99 1.00 1.00 1.00 0.93 0.94
+tel 1.00 1.00 0.99 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 0.97
+urd 1.00 0.98 0.93 0.93 0.97 0.97 1.00 0.99 0.96 0.96 0.97 0.97 0.98 0.98 1.00 1.00 0.81 0.79
+Table 6: F1-Scores on ILID Dev and Test Sets for each classifier on each language. Highest scores in each language in respective datasets across models are marked in bold